OpenStudy (anonymous):

i need the diffrence quotient for f(x) = 2x + 4

OpenStudy (anonymous):

\[\frac{2(x+h)+4-2x-4}{h}=\frac{2x+2h+4-2x-4}{h}=\frac{2h}{h}=2\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw